Circular No. 7173

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 K1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet that he discovered in images taken with the C3 coronagraph aboard SOHO. On May 20.47 UT, the comet was ∼ 11 R_{⊙} from the sun and showed a tail; on May 20.51, V = 6.8. Astrometric measurements made by Biesecker (and reduced by B. G. Marsden, Smithsonian Astrophysical Observatory), together with orbital elements by Marsden, are given on MPEC 1999-K14.

1999 UT  α_{2000}  δ_{2000}
May 20.279  3^{h}54.^{m}6  +16^{o}59′

SUPERNOVA 1999ce IN ANONYMOUS GALAXY P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999ce was obtained with the 1.5-m Tillinghast telescope on May 18.2 UT that shows it to be a type-Ia supernova near maximum brightness. Narrow emission lines in the host galaxy provide a redshift of 23 400 km/s. The photospheric velocity derived from the minimum of the Si II absorption is 12 100 km/s.

SUPERNOVA 1999aq IN ANONYMOUS GALAXY Corrigendum. L.-G. Strolger reports that the coordinates for this su-

pernova on IAUC 7125 were inadvertently provided for equinox 1950.0. For equinox 2000.0, α = 9^{h}38^{m}10^{s}.79, δ = −5^{o}08′56′′.2.

SUPERNOVA 1999by IN NGC 2841 Visual magnitude estimates, contributed in part by J. A. Mattei,

AAVSO: May 4.86 UT, 13.5 (T. Kinnunen, Espoo, Finland); 5.974, 13.7 (B. H. Granslo, Fjellhamar, Norway); 6.842, 13.4 (K. Hornoch, Lelekovice, Czech Republic); 8.995, 13.2 (P. Schmeer, Bischmisheim, Germany); 9.92, 13.2 (R. J. Bouma, Groningen, The Netherlands); 11.40, 13.6 (N. Biver, Oahu, HI); 12.017, 13.6 (M. Suárez, Las Palmas, Canary Islands); 15.844, 13.3 (K. Hornoch, Lelekovice, Czech Republic); 16.90, 13.2 (J. Carvajal, Madrid, Spain); 18.899, 13.4 (M. Reszelski, Szamotuly, Poland).

1999 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green